OpenStudy (anonymous):

-48mx divided by -8x

jimthompson5910 (jim_thompson5910):

-48 divided by -8 is (-48)/(-8) = 6 mx divided by x is (mx)/x = mx^(1-1) = mx^0 = m(1) = m So the answer is 6m

OpenStudy (anonymous):

6m, -48/-8=6 mx/x is just m for example if you replace the variable with a number....like 5 so you have m5/5....5/5 is 1. so you are left with m1 which is the same as m.